OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. Question 1 options: 1) y - 6 = -3x - 3 2) y - 6 = -3(x + 1) 3) y = -3x + 3 4) 3x + y = 3

OpenStudy (lgbasallote):

the point slope form is \[y - y_1 = m(x- x_1)\] just substitute \[y - (6) = -3(x - (-1) )\] simplify that and you have your answer